OpenStudy (anonymous):

Graph the inequality on a plane. 2x+3y<6...translate to y=mx+b form.

OpenStudy (anonymous):

\[2x+3y \le6\]

OpenStudy (saifoo.khan):

\[y \le \frac{-2y+6}{3}\]

OpenStudy (saifoo.khan):

\[y \le \frac{-2y}{3} +2\]